### v2

# Upgrade of the PandaX-4T online data quality monitoring system and perspectives for future multi-tons PandaX upgrades

Abstract
PandaX-4T is a xenon-based multi-purpose experiment, focusing on particle and astrophysics research. The data quality monitoring system plays a crucial role in the experiment. This system enables the prompt detection of potential issues during data collection. In order to meet the upgrade requirements of the experiment, we have implemented several updates to improve overall data throughput and provide users with more comprehensive information. As a result, the system is capable of monitoring half of the collected data in future operations of the PandaX-4T experiment. Furthermore, with updated hardware, the system is also well-equipped to meet the requirements of the future multi-ten-tonne level PandaX-xT experiment.
1 Introduction
The PandaX-4T experiment [1] is a multi-tonne liquid xenon observatory located in the China Jinping Underground Laboratory  [2]. It utilizes the dual-phase liquid xenon technology [3] to explore a wide range of physics phenomena, such as dark matter detection [4], coherent elastic scattering of solar neutrinos [5, 6], double beta decay of xenon isotopes [7] and other rare events [8]. After completing the initial data taking stage (Run1), a detector upgrade was carried out. The recent successful completion of the detector upgrade has paved the way for the scheduling of new data taking stage (Run2). The future plans include the development of PandaX-xT, a multi-ten-tonne observatory that will operate at the same site [9].
Scattering events occurring in the active region of the dual-phase xenon detector may produce prompt scintillation light signal ($S1$) in liquid xenon and electroluminescence signal ($S2$) in gas xenon. These signals can be captured by two arrays of photomultiplier tubes (PMTs) on the top and bottom of the detector, read out separately with 169 and 199 channels. The captured signals are digitized using flash analog-to-digital converters (FADCs) and subsequently stored for further analysis. Due to the combination of high digitization rate of 250 MHz (500 MHz post-upgrade [10]) and a large number of readout channels, a substantial amount of data is generated. It is estimated that PandaX-4T is collecting approximately 1 PB of data per year [11], while PandaX-xT is projected to produce over 1.8 PB of data per year [9]. These data are stored in a custom format based on the Bamboo Shoot3 library [12]. During the data taking process, an Apache Kafka message queue [13, 14, 15] is employed to transmit metadata information to various functional modules. These modules encompass the online data processing, data transfering and data quality monitoring system.
The data quality monitoring system plays a crucial role in the data taking process by offering reference plots and numerical values from the collected data. These resources serve as valuable references for both the on-site operators and the data analyzers. In conjunction with the upgrade of the current PandaX-4T experiment, the doubled sampling rate results in a higher data flow. This poses challenges for the existing data quality monitoring system to extract reference resources in a timely manner. To address this, we have performed a comprehensive upgrade of the system, making it more suitable to meet the requirements of both the upcoming operation of PandaX-4T and the future PandaX-xT experiment.
This paper provides a comprehensive description of the architecture and upgrade of the data quality monitoring system. In Sec. 2 we present an introduction to the data architecture of the system, together with its sub-modules. Sec. 3 focuses on the upgrade of the data quality information extraction, aimed at improving data throughput. In Sec. 4, we introduce an enhancement of the web-based user interface, which involves a precise presentation of data. Finally, a summary is provided in Sec. 5.
2 The data quality monitoring system
In this section, we introduce the architecture, components, and functionalities of the data quality monitoring system.
2.1 Architecture of the system
The data quality monitoring system is essential for swiftly identifying issues during data collection. It must extract physical information from processed data obtained through the online data processing system once the data is ready. Furthermore, the metadata of runs is stored in the database. Thus, the data quality monitoring system is tightly integrated with the online data processing system and the database, as illustrated in Fig. 1. The system comprises three independent sub-modules: a) the plots generator, responsible for extracting information from processed files; b) the Web UI, which serves as the user interface; c) the app server, reading extracted information from the database or file system and transmitting it to the user. The interactions among these sub-modules are distinctly represented in Fig. 1.
Figure 1: Architecture of the data quality monitoring system as well as the online data processing system. The blue lines indicate the direction of the data flow. The dashed orange lines in the diagram represent the information exchange with the database, while the solid yellow lines represent the information exchange with the message queue. The online data processing system processes specific raw data produced by the data acquisition (DAQ) program. The plots generator extracts reference quality information from the processed data like radon event rate and electron lifetime, which is then saved as plots and data files or stored in the database. Users can access this information through the WebUI, which interacts with the app server to retrieve the required files or data from the database.
The different systems within the architecture are interconnected using the Apache Kafka message queue. This message queue allows for the creation and sharing of different “topics”, which serve as categories for organizing and sharing messages among the various systems involved [13, 14, 15]. In the DAQ system, raw data from a single run are stored as a series of files, each assigned a unique file number for identification. After completing the writing process for a data file, a message containing the run and file number is sent to the “daq” topic. The online data processing system consumes this message and determines whether or not to process the raw data based on predefined rules. The data processing steps employed by the system follow the standard procedures introduced in Ref. [11]. Once the processing of a given file is completed, the same message is sent to the “dq”(data quality) topic to notify the plots generator about extracting the data quality information. In the previous version of the plots generator, about 50 image files were generated for different variables that are valuable for quality monitoring, and some information is stored directly to the database.
To access the plots and related information, users can utilize their browsers to visit the WebUI, which is served by a Nginx web server  [16]. The WebUI, developed using the React library [17], initially displays the run list. When a user selects a specific run from the interface, the corresponding data quality plots and values are fetched from the back-end app server and displayed below the run list. The example screenshot is shown in Fig. 2.
Figure 2: Screenshot of the interface displaying run-related data quality information. (Note: Image is modified to fit the page. Grayed-out sections have been omitted.)
The back-end app server is built using the lightweight Echo framework [18]. It offers representational state transfer (REST) application programming interfaces (APIs) that allow users to access the related data quality information.
2.2 Plots generator
For each selected raw data file, the online data processing system generates three output files, containing hit, signal, and event level information, respectively. A hit corresponds to a pulse in the raw waveform of a PMT, containing information about detected photons. A signal is a grouping of hits from multiple PMTs with similar timing. An event consists of signals occurring within a specific time window, potentially originating from the same physical interaction [11]. Once the processing of a data file is completed, the plots generator is notified through the message queue. The multi-threaded plots generator operates with one thread dedicated to receiving messages and placing them into an internal data structure, which is a mapping between run numbers and the list of corresponding file numbers. Upon the arrival of a message for a new run, a dedicated thread is assigned to generate data quality plots for that particular run. The thread analyzes the output files in the order specified in the corresponding file number lists. Once all the related files have been analyzed, the file number is removed from the list, and the files are subsequently removed from storage. The thread is recycled for a new run when the file list remains empty for a minimum of one hour.
Presently, more than 50 distributions about data quality variables are derived from the processed data. From the files containing hit level information, variables such as hit rate, noise-to-signal pulse ratio, stability of channel baseline, and hit areas are extracted. The status of PMTs is monitored using the collected charges of each PMT based on the signal data. Additionally, the rate of small $S1$ signals and sparking signals is monitored to aid identifying potential detector malfunctions. Various signal properties are collected for distribution analysis. At the event level, the lightened PMTs of events with single $S1$ or single $S2$ signals are monitored, along with other types of special events. For the “good” events that consist of one major $S1$ followed by one major $S2$, the top-bottom asymmetry of the $S1$ signal and the time separation between S1 and S2 signal serve as references for evaluating the detector status. Several variables, such as the rates of radon events and low energy electron recoil events, and the electron lifetime are calculated.
In the previous version of the plots generator, the distributions were stored in memory using TH1D or TH2D objects from the ROOT [19] framework. These objects were updated as the input files were analyzed. The plots were regenerated in PNG format with updated information after processing every 10 raw files. The calculated variables were simultaneously uploaded to the database.
2.3 WebUI
The WebUI presents information in six categories across different pages, including runs, files, channels, rates, PMT gains, and durations. Most of the data quality information is specific to each run. By default, the interface exhibits a list of the 50 latest runs. Above this list, searching boxes are provided to enable users to search for specific runs based on the run number, run number range, keywords in the description, or the type. Each entry in the list corresponds to a run and can be edited by the shifter to update the run type and description. Upon clicking on a run entry, the data quality plots and numbers corresponding to that particular run will be displayed below the run list.
The page for displaying file information primarily serves the purpose of manually tagging abnormal or bad files. On the other hand, the page dedicated to displaying channel information is utilized for tagging bad channels. The rates page offers a comprehensive view of the evolution of key variables over time, including the electron lifetime, hit rates, signal rates, and rates of specific events. These variables serve as references for assessing the long-term stability of the detector. The gains page allows users to examine the gain evolution of each PMT and explore the gain pattern for specific runs. The duration page offers a comprehensive view of the accumulated run duration of different types of data. It displays the duration of the user-input run list, excluding any files that have been tagged as bad. Within these pages, the data are visualized in the browser using the EChartsJS package [20].
2.4 App server
The app server has the responsibility of reading or updating the database, searching for related files in the storage, and sending the results back to the user through the reverse proxy provided by the Nginx HTTP server. The Go-Pg library is utilized for accessing the PostgreSQL database. As the system is designed solely for internal use, there are no authentication and logging functionalities implemented in the app server. All the functionalities of the app server are exposed as REST APIs. The APIs are summarized in Table  1.
API method description
/runs GET get list of runs
/run/image/:id GET get images of a run with given id
/run/run_rates/:id GET get rates of a run with given id
/run/trigger_rates/:id GET get trigger rates of a run with given id
/run/board_rates/:id GET get board rates of a run with given id
/run/gain/:id GET get PMT gains of a run with given id
/runs/duration/type GET get duration of run of given types
/runs/duration/sum GET get duration of run within a given list
/run/:id PUT update run description and type
/run/channels PUT set bad channels
/files GET get list of files
/files/update PUT update file quality information
Table 1: The APIs provided by the app server.
3 Upgrade of the plots generator
The updated plots generator offers a significant increase in throughput compared to the initial version. The updated plots generator also includes support for the prompt acquisition of self-corrected PMT gain by utilizing online generated hit data.
3.1 Performance optimization
The first version of plots generator was designed to handle input data sequentially. It has been effective in processing data with a low data rate, such as background data, by processing one file out of every two files. However, during a high bandwidth calibration run, later files experience longer waiting times for processing. To avoid the blocking, the fraction of processed files was reduced to 1/8 in Run0 and Run1 calibration process, but it may result in the generated plots being less representative of the collected data.
Another concern with the current plots generator is the generation of images using the “Print()” function of TH1D or TH2D classes from ROOT, which resulting in performance downgrade. On the server equipped with two AMD EPYC 7452 CPUs and 256 GB of memory, the operation to generate all the required data quality images took approximately 3.76 seconds on average. Additionally, as the number of data runs increases, the accumulation of small image files becomes a burden on the file system and difficult to manage.
In the upcoming upgrade of PandaX-4T and future PandaX-xT experiments, the data rates are expected to be significantly higher. To meet the new demand based on the previously mentioned drawbacks, two major updates have been implemented.
At first, we made an update by replacing the internal data structure to store data quality information from TH1D and TH2D to vectors in the C++ standard library. To improve efficiency, we have decided not to generate image files. Instead, we now dump the data quality data of a run into a single text file in JSON format. For each run, the text file that contains all the data quality information has a size of approximately 7 MB.
Another significant optimization involves utilizing all available threads in the pre-allocated thread pool to concurrently process incoming files. This allows files with different file numbers to be processed in parallel, improving overall efficiency. Before updating the vectors used for storing data quality information, the working thread must acquire the corresponding lock to prevent data racing.
The overall improvement in processing speed is studied with the PandaX-4T data from various types of runs. To check the performance, a thread pool with only one thread is used, as in the previous implementation of the program. The results are presented in Tab. 2. It can be observed that the throughput has improved by over 50% with only one working thread.
Run type processing time (minutes) throughput (MB/s) improvement
previous updated previous updated in throughput
background 194.15 119.37 19.34 31.45 62.6%
Rn 201.43 123.93 18.64 30.30 64.3%
83mKr 191.35 118.73 19.62 31.62 61.2%
DD 213.13 130.42 17.62 28.79 63.4%
241AmBe 203.28 132.25 18.47 28.39 53.7%
232Th 209.1 132.00 17.96 28.44 58.4%
60Co 215.23 135.17 17.44 27.78 59.2%
Table 2: Comparison of processing time and throughput for two versions of plots generators on various types of data. For each run, the initial 200 files are processed. Each file has a size of approximately 1.1 GB. The updated version utilizes only one thread for data processing.
By increasing the number of threads in the pre-allocated thread pool, the throughput of the updated plots generator exhibits nearly linear improvement initially. However, after surpassing 10 threads, the improvement becomes less pronounced, as depicted in Fig. 3. This behavior can be attributed to the constraints imposed by concurrent reading from the storage and the need to maintain data integrity through locking mechanisms in a multi-threaded environment.
Given the constraint of limited available CPU cores on the server, a careful consideration was made to strike a balance between data processing and plot generation. Ultimately, a decision was made to use 4 threads in the pool. This configuration allows for efficient data monitoring while simultaneously accommodating online data processing. With this setup, the plots generator achieves a final throughput of 100-120 MB/s. Considering that the data based on the Bamboo Shoot3 library is already compressed, resulting in approximately 45% space savings, this throughput can be translated to around 180-220 MB/s for the raw binary data from the DAQ. Based on previous experience with the operation of PandaX-4T, the raw data rate for background runs is approximately 30 MB/s, while for Rn calibration runs it is around 160 MB/s. With the new DAQ system’s doubled sampling rate, the raw data rate will double as well, reaching 320 MB/s. Selecting 4 threads enables the online analysis of half of the generated data, providing valuable insights into the data quality. To meet the demands of future experiments, upgrading the server with additional CPU cores would allow for more threads, and consequently a higher throughput, addressing the need for increased data processing capacity.
Figure 3: Throughput (MB/s) of the plots generator with allocated number of threads for different types of data. The dashed blue line indicates the linear reference.
3.2 Support for the extraction of the PMT gain
The PandaX-4T experiment utilizes weekly calibration with light-emitting diodes (LEDs) to calculate PMT gains. These LEDs, driven by a standard driver circuitry with adjusted voltage, can generate fast and faint light pulses with appropriate occupancies [21, 22]. This allows for the clean separation of single photoelectrons (SPEs) which are used to fit the PMT gains, also known as LED gains. However, this method may encounter challenges such as uneven illumination and insufficient SPE samples [23]. To address these issues, a recent analysis of PandaX-4T [24] adopts SPEs from physics data to correct the PMT gains, known as self-corrected gains, based on the calculated LED gains.
The self-corrected gains of the PMTs were extracted from every 1 in 100 files during a massive data reprocessing. This was accomplished by identifying the single photon hits within the selected data, and fitting their charge distribution. The charge values used for the fitting were obtained with the LED gains. In an ideal case, it was expected that the fitted peak with Gaussian function would be at 1 photoelectron (PE). The fitted value is used to correct the PMT gain. The corrected values are stored in the database and utilized as self-corrected gains for the hit charges during the subsequent analyses.
The inclusion of self-corrected gains extraction in the offline data processing chain resulted in a 17% increase in processing time for the complete processing of one raw file. This increase is due to the need to obtain information from a newly generated file with hit information, which is skipped in the default offline data processing. In contrast, the hit file is generated by default in online processing, as it is required by the plots generator to calculate hit-related variables. The update adding the functionality of extracting self-corrected gain values to the plots generator did not result in a noticeable increase in processing time. The benefits of this update extend beyond time-saving measures. By leveraging the plots generator, lower statistical errors in the determination of self-corrected gains can be achieved compared to the previous methods, as more data files will be processed. Additionally, this update aids in the identification of potential issues with PMTs during data collection.
4 Upgrade of the web application
Incorporating JSON-based quality information requires an update to the WebUI to enable the direct visualization of this data in the browser. The JSON file organizes different data categories as an array of mappings, each of which corresponds to a specific name and its associated data. The WebUI utilizes the EChartJS package to generate plots for each element in the array. The current WebUI predominantly employs three types of plots: bar charts, heatmaps, and scatter plots.
Bar charts are utilized as replacements for TH1D histograms, particularly for channel-based data quality information. Heatmaps serve as substitutes for TH2D histograms, allowing for the visualization of two-dimensional distributions. Scatter plots are primarily employed to display PMT-related information. Examples of these different plot types are depicted in Figure 4. These plots can be easily exported as SVG files. Additionally, a new function has been added to the scatter plots, allowing users to view the detailed information of each data point directly.
Figure 4: Examples of bar chart(4(a)), heatmap(4(b)) and scatter plot(4(c)).
Plotting multiple images simultaneously in the browser can lead to significant performance issues. To address this, we have used the lazy plotting technique, with which the plotting of images is deferred until they are within the browser window, ensuring efficient resource utilization.
To ensure the continuous monitoring of data quality information during an ongoing data collection run, a new auto-refreshing function has been implemented. This function automatically requests new data every 30 seconds and updates the related plots accordingly. This feature proves to be extremely valuable for shifters, as it allows them to stay up-to-date with the latest data quality information.
Minor updates have been made to the back-end app server as well, including the addition of two REST APIs. These APIs enable users to obtain the JSON information of a run and retrieve the self-corrected gain information.
In addition to the aforementioned modifications, the configuration of the Nginx web server has been updated to enable gzip compression for text data, which automatically compresses the JSON file during network transfer. As a result, a 7 MB JSON file is compressed to approximately 200 KB, significantly reducing network bandwidth consumption.
5 Summary
The data quality monitoring system plays a crucial role in ensuring prompt detection of potential data-related issues during the PandaX-4T data collection. In this paper, we have presented the framework of the data quality monitoring system in the PandaX-4T experiment, and most importantly the upgrades made to its critical components.
Regarding the plots generator, we have made several updates. Firstly, we have revamped the internal data structure by replacing the histogram classes from ROOT with standard containers. Secondly, we have implemented multi-threading processing of incoming files, incorporating a locking mechanism to maintain data integrity during concurrent updating of the run level information. Additionally, we have transitioned from using image files to using JSON files for extracting data quality information. These updates have significantly improved the overall data throughput, allowing us to monitor half of the collected data online in the future operations of the PandaX-4T experiment.
As for the WebUI, our update focuses on client-side plotting based on JSON data obtained from the app server. This update reduces the requirement of network bandwidth and provides users with more extensive information compared to the previous implementation.
The updated data quality monitoring system successfully fulfills all necessary requirements for the upcoming operation of the PandaX-4T experiment. Furthermore, it demonstrates the potential to effectively monitor data with higher bandwidth in the future PandaX-xT experiment, especially when running on a more powerful server.
Acknowledgments
This project is supported by the grants from National Natural Science Foundation of China (No. 12175139).
Conflict of Interest Statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.